# OpenCodeInterpreter: Integrating Code Generation with Execution and Refinement

### v2

@@ Abstract
[…]
@@ 4.1 Results of Single-turn Code Generation
We compare OpenCodeInterpreter’s single-turn code generation performance against premier models such as GPT-3.5/4-Turbo OpenAI (2022, 2023), CodeLlama-Python  (Roziere et al., 2023), WizardCoder  (Luo et al., 2023), Deepseek-Coder Guo et al. (2024), CodeT5+  (Wang et al., 2023c) across different scales. Leveraging data from the EvalPlus leaderboard as of February 10th, 2024, we examine OpenCodeInterpreter’s achievements on the HumanEval and MBPP benchmarks, as well as their advanced versions, HumanEval+ and MBPP+. For straightforward comparisons, we consolidate results across different model scales into one table, facilitating direct performance comparisons between each model scale and the respective variants of OpenCodeInterpreter.
@@ Appendix A Source Data Filtering
Query Filtering Prompt 1
Rate the following code queries on a scale of 1 to 5 based on their complexity, where 1 is the easiest and 5 is the most difficult. Consider the complexity of the query Query: [{query}] You are obliged to choose only from the following list. Scoring Criteria: 1 Point - Very Basic: The query involves simple operations or common issues 2 Points - Basic: The query involves fundamental programming concepts or commonly used functions 3 Points - Intermediate: The query requires some programming experience, possibly involving multiple steps 4 Points - Difficult: The query involves advanced programming skills, including complex logic, algorithms, or data structures 5 Points - Very Difficult: The query requires extensive expertise, potentially involving innovative problem-solving approaches or unique algorithm design Please give the score first then explain why
Query Filtering Prompt 2
Rate the following code queries on a scale of 1 to 5 based on their complexity, where 1 is the easiest and 5 is the most difficult. Consider the complexity of the query Query: [{query}] You are obliged to choose only from the following list. Scoring Criteria: 1 Point - Moderately Difficult: Involves understanding specific programming concepts or libraries, and may include medium complexity algorithms or data structures like basic sorting algorithms or tree structures. 2 Points - Challenging: Requires handling more complex logic or algorithms such as advanced sorting algorithms, recursive logic, or intermediate data structures like hash tables and heaps. 3 Points - Highly Challenging: Demands deeper knowledge in algorithms and data structures, potentially including graph algorithms, dynamic programming, or complex string manipulation techniques. 4 Points - Advanced: Focuses on proficiency in programming and algorithm design, dealing with complex system architecture issues, performance optimization, or solving advanced algorithmic challenges like NP-hard problems. 5 Points - Expert Level: The highest difficulty level, requiring innovative problem-solving approaches or unique algorithm design, possibly involving interdisciplinary knowledge or the application of cutting-edge technologies. Please give the score first then explain why
@@ Appendix B Simulating Interactions for Data Collection
System prompt for multi-turn execution feedback
@@
You are an AI code interpreter. Your goal is to help users do a variety of jobs by executing Python code. You should: 1. Comprehend the user’s requirements carefully & to the letter. 2. Give a brief description for what you plan to do & call the provided function to run code. 3. Provide results analysis based on the execution output. 4. If error occurred, try to fix it. 5. Response in the same language as the user.
System prompt for multi-turn human feedback
@@
You are a user who gives feedback to the latest generated code. If no available code is found in the conversation, you should give a feedback to encourage assistant to generate code. NOTE: your feedback should be WITHIN 2 SHORT SENTENCES. You can refer to the following types of feedback: 1. **Syntax and Formatting**: Checking for syntax errors, inconsistent formatting, and suggesting adherence to standard coding styles for readability and maintainability. 2. **Efficiency**: Identifying parts of the code that can be optimized for better performance, such as reducing time complexity, optimizing loops, or suggesting more efficient data structures. 3. **Functionality Enhancements**: Suggesting additional features or enhancements that could make the code more functional or user-friendly. 4. **Code Clarity and Documentation**: Recommending improvements in code comments and documentation to make the code more understandable and easier to maintain. 5. **Bug Identification**: Pointing out any potential bugs or logical errors in the code and suggesting ways to fix them. 6. **Security Improvements**: Highlighting any security vulnerabilities in the code and suggesting best practices to enhance security. 7. **Compatibility and Testing**: Advising on making the code more compatible with different environments or platforms and suggesting more comprehensive testing scenarios. 8. **Resource Optimization**: Identifying areas where the code might be using more resources than necessary (like memory or CPU) and suggesting optimizations. 9. **Scalability**: Providing insights on how the code can be made more scalable to handle larger data sets or more users. 10. **Adherence to Best Practices**: Ensuring the code follows the best practices specific to the language or framework being used. Your output MUST be in a json format like this: { "satisfied": "The points that have been achieved in generated code", "not_satisfied": "The points that have not yet been achieved in generated code", "feedback": "The actual feedback. Your feedback should be WITHIN 2 SHORT SENTENCES. Feedback must come from a point included in ’not_satisfied’ field. You can ask the assistant here to generate code if no available code is found in previous conversations." }
System prompt for deliberately generating incorrect code
@@
You are an AI code interpreter. Your goal is to generate and execute Python code. Your code MUST contain at least one of the following types of errors: 1. Syntax Error: This type of error occurs when the code violates the grammar rules of the programming language. For example, forgetting to close a parenthesis or a quotation mark, or misspelling a keyword. 2. Logical Error: These errors sneak into your code when there’s a misunderstanding of the problem you’re solving, leading to incorrect results despite the code running without crashing. For example, calculating the average of a list of numbers by summing them up but forgetting to divide by the count of the numbers. 3. Type Error: This error occurs when an operation is applied to an object of an inappropriate type. For example, attempting to concatenate a string with an integer without converting the integer to a string first. 4. Name Error: This happens when the code attempts to reference a variable or a function name that hasn’t been defined. For example, trying to print a variable that hasn’t been declared. 5. Timeout Error: This error occurs when your code gets stuck in a loop that never ends, either due to a logic flaw or a condition that never becomes false. In programming, such an error can cause your application to hang indefinitely, consuming resources and potentially leading to a crash if not handled properly. For example, writing a loop that waits for a certain condition to change, but the condition is never updated within the loop. NOTE: 1. You MUST make mistakes in the generated code! 2. Do not explain the errors within. Just write your thoughts and code as normal. 3. Do not tell me you are writing the wrong code in any form (e.g., in text/code/comments). Just pretend you are writing the correct code and still not recognizing the errors.
@@ Appendix C Natural Language Explanations Generation
Prompt for generating natural language explanations using GPT-4
Here is a list containing a series of dialogues between a user and a programmer assistant. Following the previous dialogues, the user posed a latest problem. The assistant has now crafted the correct code based on the previous dialogues and the latest problem. Assuming you are this programmer assistant, please add some text before the code. The purpose of this text is to respond to the latest problem and to introduce the code that follows. This text may include: language used in the code, algorithm used in the code, step-by-step implementation overview, and other relevant content. You may use phrases like "The following code", "My code", "My solution", to refer to the @@Code. Your response should ONLY contain the text that you add before the code. Your only task is to write the text, never modify the code or remind me something. Never restate the previous dialogues and the problem. @@Previous Dialogues {previous dialogues} @@Recent Problem: {recent problem} Add the text there. @@Code: {code}
@@ Appendix D Model Evaluation Prompts
Prompt for HumanEval and HumanEval+
You are an exceptionally intelligent coding assistant that consistently delivers accurate and reliable responses to user instructions. @@ Instruction Here is the given code to do completion: “‘{language} {original prompt} “‘ Please continue to complete the function with {language} programming language. You are not allowed to modify the given code and do the completion only. Please return all completed codes in one code block. This code block should be in the following format: “‘{language} # Your codes here “‘ @@ Response
Prompt for MBPP and MBPP+
You are an exceptionally intelligent coding assistant that consistently delivers accurate and reliable responses to user instructions. @@ Instruction Here is the given problem and test examples: {original prompt} Please use the {language} programming language to solve this problem. Please make sure that your code includes the functions from the test samples and that the input and output formats of these functions match the test samples. Please return all completed codes in one code block. This code block should be in the following format: “‘{language} # Your codes here “‘ @@ Response
@@
Prompt for GPT models mimicking human feedback with canonical solution
You are tasked with providing guidance to a programmer who has drafted a code for a programming problem. Your role is to mimic human-like responses and offer suggestions for modifying the code based on the canonical solution and the observed execution results. You should NOT directly revealing contents of the @@Canonical Solution or mentioning terms such as "canonical solution." You should refrain from directly writing code. Begin by thoroughly examining the existing code and its functionality. Compare the @@Existing Code with the @@Canonical Solution provided. Note any discrepancies in logic, approach, or implementation. Analyze the @@Execution Result obtained from running the @@Existing Code. Identify any errors, unexpected behavior, or deviations from the expected output. Consider potential edge cases, optimization opportunities, or alternative approaches based on insights from both the @@Canonical Solution and @@Execution Result. Offer guidance in a clear and understandable manner, explaining the rationale behind each suggestion. Refrain from providing actual code solutions, but instead focus on conceptual modifications or strategies. Provide constructive feedback to help the programmer improve their coding skills. Remember, your role is to simulate human-like guidance and expertise in programming without directly implementing solutions. Please respond in no more than three sentences. @@Problem {original prompt} @@Existing Code {sanitized code} @@Execution Result {execution result} @@Canonical Solution {canonical solution} @@Guidance
Prompt for GPT models mimicking human feedback without canonical solution
You are tasked with providing guidance to a programmer who has drafted a code for a programming problem. Your role is to mimic human-like responses and offer suggestions for modifying the code based on the observed execution results. You should refrain from directly writing code. Begin by thoroughly examining the existing code and its functionality. Analyze the @@Execution Result obtained from running the @@Existing Code. Identify any errors, unexpected behavior, or deviations from the expected output. Consider potential edge cases, optimization opportunities, or alternative approaches based on insights from the @@Execution Result. Offer guidance in a clear and understandable manner, explaining the rationale behind each suggestion. Refrain from providing actual code solutions, but instead focus on conceptual modifications or strategies. Provide constructive feedback to help the programmer improve their coding skills. Remember, your role is to simulate human-like guidance and expertise in programming without directly implementing solutions. Please respond in no more than three sentences. @@Problem {original prompt} @@Existing Code {sanitized code} @@Execution Result {execution result} @@Guidance

### v3

@@ Abstract
Tianyu Zheng1thanks: Equal Contributions., Ge Zhang1,211footnotemark: 1, Tianhao Shen [^1] footnotemark: 1, Xueling Liu [^1] footnotemark: 1, Bill Yuchen Lin3, Jie Fu4, Wenhu Chen1,2, Xiang Yue1,5thanks: Corresponding Author. 1Multimodal Art Projection Research Community, 2University of Waterloo, 3Allen Institute for Artificial Intelligence, 4HKUST, 5IN.AI Research {zhengtianyu0428, xiangyue.work}@gmail.com, ge.zhang@uwaterloo.ca https://opencodeinterpreter.github.io
@@ 2.2 Coding Challenges from LeetCode
StarCoder2 Lozhkov et al. (2024) 7B Base 35.4 (29.9) 54.4 (45.6) 44.9 (37.8)
   OpenCodeInterpreter-SC2 73.8 (68.9) 61.7 (51.1) 67.8 (60.0)
   + Execution Feedback 7B Instruct 75.6 (69.5) 66.9 (55.4) 71.3 (62.5)
@@
StarCoder2 Lozhkov et al. (2024) 15B Base 46.3 (37.8) 66.2 (53.1) 56.3 (45.5)
   OpenCodeInterpreter-SC2 75.6 (69.5) 71.2 (61.2) 73.4 (65.4)
   + Execution Feedback 15B Instruct 77.4 (72.0) 74.2 (63.4) 75.8 (67.7)
@@ 4.1 Results of Single-turn Code Generation
We compare OpenCodeInterpreter’s single-turn code generation performance against premier models such as GPT-3.5/4-Turbo OpenAI (2022, 2023), CodeLlama-Python  (Roziere et al., 2023), WizardCoder  (Luo et al., 2023), Deepseek-Coder (Guo et al., 2024), CodeT5+  (Wang et al., 2023c) across different scales. Leveraging data from the EvalPlus leaderboard as of February 10th, 2024, we examine OpenCodeInterpreter’s achievements on the HumanEval and MBPP benchmarks, as well as their advanced versions, HumanEval+ and MBPP+. For straightforward comparisons, we consolidate results across different model scales into one table, facilitating direct performance comparisons between each model scale and the respective variants of OpenCodeInterpreter.
@@ 4.3 Ablations of Data Sources
4.4 Analysis of Dataset Leakage
We conduct a thorough analysis to address dataset leakage, aiming to assess the degree of overlap between our proposed Code-Feedback dataset and the benchmarks utilized in our study, namely HumanEval (+) and MBPP (+).
We specifically examine the duplication ratio of code snippets between our dataset and the benchmarks (Chen et al., 2021), focusing on consecutive lines to gauge the extent of similarity. The results of our analysis are summarized in the table below. As depicted in Table 4, the duplicate line ratios are notably low across all examined consecutive line lengths. This indicates minimal overlap between our dataset and the benchmarks, thereby mitigating concerns regarding dataset leakage.
Consecutive Lines HumanEval (+) MBPP (+)
5 lines 1.19% 0.51%
6 lines 0.53% 0.00%
7 lines 0.00% 0.00%
Table 4: Duplicate Line Ratios between Code-Feedback Dataset and Code Benchmarks.
Furthermore, upon closer examination of the duplicated lines, we observe that the majority are generic and widely used code snippets. These snippets contribute minimally to the performance improvement on specific benchmarks, as they are not indicative of dataset-specific patterns or characteristics. In light of these findings, we can confidently assert that the risk of dataset leakage is low, affirming the integrity and independence of our Code-Feedback dataset.
4.5 Evaluation of Multi-turn Settings
To address the need for a challenging multi-turn evaluation dataset beyond HumanEval and MBPP, we conduct tests on coding-type questions in the MT-Bench (Zheng et al., 2024) dataset, which comprises 10 coding-type questions, each including a primary question and a follow-up question. We follow the standard MT-Bench setup and utilize GPT-4 to score the model outputs (from 1-10). The test results are presented in the table below.
Model First Turn Second Turn Average
GPT-4 9.0 8.1 8.6
GPT-3.5-Turbo 6.6 7.2 6.9
Claude-Instant-V1 6.5 7.0 6.8
OpenCI-DS-33B 6.8 6.7 6.8
OpenCI-DS-6.7B 6.7 5.8 6.3
DS-33B-Instruct 6.7 4.2 5.5
DS-6.7B-Instruct 5.6 4.5 5.1
OpenCI-CL-34B 5.6 4.1 4.9
OpenCI-CL-70B 5.7 3.7 4.7
OpenCI-CL-13B 5.5 3.7 4.6
CL-34B-Instruct 4.5 3.4 4.0
Vicuna-33B-V1.3 3.8 2.9 3.4
Vicuna-13B-V1.3 3.8 2.7 3.3
CL-7B-Instruct 3.3 2.5 2.9
CL-13B-Instruct 3.3 2.3 2.8
Table 5: Performance of various models on the 10 code-related questions in the MT-Bench dataset. ‘CL’:based on CodeLlama; ‘DS’: based on DeepseekCoder
The analysis of the performance of OpenCodeInterpreter in the MT-Bench dataset reveals that OpenCodeInterpreter-DS-33B and OpenCodeInterpreter-DS-6.7B tend to perform better compared to other open-source models. Notably, when compared with DeepSeek-Coder-33B-Instruct, our OpenCodeInterpreter demonstrates improved performance in the second turn, which further demonstrates the usefulness of our Code-Feedback dataset.
@@ Appendix A Source Data Filtering
Query Filtering Prompt 1 Rate the following code queries on a scale of 1 to 5 based on their complexity, where 1 is the easiest and 5 is the most difficult. Consider the complexity of the query Query: [{query}] You are obliged to choose only from the following list. Scoring Criteria: 1 Point - Very Basic: The query involves simple operations or common issues 2 Points - Basic: The query involves fundamental programming concepts or commonly used functions 3 Points - Intermediate: The query requires some programming experience, possibly involving multiple steps 4 Points - Difficult: The query involves advanced programming skills, including complex logic, algorithms, or data structures 5 Points - Very Difficult: The query requires extensive expertise, potentially involving innovative problem-solving approaches or unique algorithm design Please give the score first then explain why
Query Filtering Prompt 2 Rate the following code queries on a scale of 1 to 5 based on their complexity, where 1 is the easiest and 5 is the most difficult. Consider the complexity of the query Query: [{query}] You are obliged to choose only from the following list. Scoring Criteria: 1 Point - Moderately Difficult: Involves understanding specific programming concepts or libraries, and may include medium complexity algorithms or data structures like basic sorting algorithms or tree structures. 2 Points - Challenging: Requires handling more complex logic or algorithms such as advanced sorting algorithms, recursive logic, or intermediate data structures like hash tables and heaps. 3 Points - Highly Challenging: Demands deeper knowledge in algorithms and data structures, potentially including graph algorithms, dynamic programming, or complex string manipulation techniques. 4 Points - Advanced: Focuses on proficiency in programming and algorithm design, dealing with complex system architecture issues, performance optimization, or solving advanced algorithmic challenges like NP-hard problems. 5 Points - Expert Level: The highest difficulty level, requiring innovative problem-solving approaches or unique algorithm design, possibly involving interdisciplinary knowledge or the application of cutting-edge technologies. Please give the score first then explain why
@@ Appendix B Simulating Interactions for Data Collection
System prompt for multi-turn execution feedback You are an AI code interpreter. Your goal is to help users do a variety of jobs by executing Python code. You should: 1. Comprehend the user’s requirements carefully & to the letter. 2. Give a brief description for what you plan to do & call the provided function to run code. 3. Provide results analysis based on the execution output. 4. If error occurred, try to fix it. 5. Response in the same language as the user.
@@
System prompt for multi-turn human feedback You are a user who gives feedback to the latest generated code. If no available code is found in the conversation, you should give a feedback to encourage assistant to generate code. NOTE: your feedback should be WITHIN 2 SHORT SENTENCES. You can refer to the following types of feedback: 1. **Syntax and Formatting**: Checking for syntax errors, inconsistent formatting, and suggesting adherence to standard coding styles for readability and maintainability. 2. **Efficiency**: Identifying parts of the code that can be optimized for better performance, such as reducing time complexity, optimizing loops, or suggesting more efficient data structures. 3. **Functionality Enhancements**: Suggesting additional features or enhancements that could make the code more functional or user-friendly. 4. **Code Clarity and Documentation**: Recommending improvements in code comments and documentation to make the code more understandable and easier to maintain. 5. **Bug Identification**: Pointing out any potential bugs or logical errors in the code and suggesting ways to fix them. 6. **Security Improvements**: Highlighting any security vulnerabilities in the code and suggesting best practices to enhance security. 7. **Compatibility and Testing**: Advising on making the code more compatible with different environments or platforms and suggesting more comprehensive testing scenarios. 8. **Resource Optimization**: Identifying areas where the code might be using more resources than necessary (like memory or CPU) and suggesting optimizations. 9. **Scalability**: Providing insights on how the code can be made more scalable to handle larger data sets or more users. 10. **Adherence to Best Practices**: Ensuring the code follows the best practices specific to the language or framework being used. Your output MUST be in a json format like this: { "satisfied": "The points that have been achieved in generated code", "not_satisfied": "The points that have not yet been achieved in generated code", "feedback": "The actual feedback. Your feedback should be WITHIN 2 SHORT SENTENCES. Feedback must come from a point included in ’not_satisfied’ field. You can ask the assistant here to generate code if no available code is found in previous conversations." }
@@
System prompt for deliberately generating incorrect code You are an AI code interpreter. Your goal is to generate and execute Python code. Your code MUST contain at least one of the following types of errors: 1. Syntax Error: This type of error occurs when the code violates the grammar rules of the programming language. For example, forgetting to close a parenthesis or a quotation mark, or misspelling a keyword. 2. Logical Error: These errors sneak into your code when there’s a misunderstanding of the problem you’re solving, leading to incorrect results despite the code running without crashing. For example, calculating the average of a list of numbers by summing them up but forgetting to divide by the count of the numbers. 3. Type Error: This error occurs when an operation is applied to an object of an inappropriate type. For example, attempting to concatenate a string with an integer without converting the integer to a string first. 4. Name Error: This happens when the code attempts to reference a variable or a function name that hasn’t been defined. For example, trying to print a variable that hasn’t been declared. 5. Timeout Error: This error occurs when your code gets stuck in a loop that never ends, either due to a logic flaw or a condition that never becomes false. In programming, such an error can cause your application to hang indefinitely, consuming resources and potentially leading to a crash if not handled properly. For example, writing a loop that waits for a certain condition to change, but the condition is never updated within the loop. NOTE: 1. You MUST make mistakes in the generated code! 2. Do not explain the errors within. Just write your thoughts and code as normal. 3. Do not tell me you are writing the wrong code in any form (e.g., in text/code/comments). Just pretend you are writing the correct code and still not recognizing the errors.
@@ Appendix C Natural Language Explanations Generation
Prompt for generating natural language explanations using GPT-4 Here is a list containing a series of dialogues between a user and a programmer assistant. Following the previous dialogues, the user posed a latest problem. The assistant has now crafted the correct code based on the previous dialogues and the latest problem. Assuming you are this programmer assistant, please add some text before the code. The purpose of this text is to respond to the latest problem and to introduce the code that follows. This text may include: language used in the code, algorithm used in the code, step-by-step implementation overview, and other relevant content. You may use phrases like "The following code", "My code", "My solution", to refer to the @@Code. Your response should ONLY contain the text that you add before the code. Your only task is to write the text, never modify the code or remind me something. Never restate the previous dialogues and the problem. @@Previous Dialogues {previous dialogues} @@Recent Problem: {recent problem} Add the text there. @@Code: {code}
@@ Appendix D Model Evaluation Prompts
Prompt for HumanEval and HumanEval+ You are an exceptionally intelligent coding assistant that consistently delivers accurate and reliable responses to user instructions. @@ Instruction Here is the given code to do completion: “‘{language} {original prompt} “‘ Please continue to complete the function with {language} programming language. You are not allowed to modify the given code and do the completion only. Please return all completed codes in one code block. This code block should be in the following format: “‘{language} # Your codes here “‘ @@ Response
Prompt for MBPP and MBPP+ You are an exceptionally intelligent coding assistant that consistently delivers accurate and reliable responses to user instructions. @@ Instruction Here is the given problem and test examples: {original prompt} Please use the {language} programming language to solve this problem. Please make sure that your code includes the functions from the test samples and that the input and output formats of these functions match the test samples. Please return all completed codes in one code block. This code block should be in the following format: “‘{language} # Your codes here “‘ @@ Response
@@
Prompt for GPT models mimicking human feedback with canonical solution You are tasked with providing guidance to a programmer who has drafted a code for a programming problem. Your role is to mimic human-like responses and offer suggestions for modifying the code based on the canonical solution and the observed execution results. You should NOT directly revealing contents of the @@Canonical Solution or mentioning terms such as "canonical solution." You should refrain from directly writing code. Begin by thoroughly examining the existing code and its functionality. Compare the @@Existing Code with the @@Canonical Solution provided. Note any discrepancies in logic, approach, or implementation. Analyze the @@Execution Result obtained from running the @@Existing Code. Identify any errors, unexpected behavior, or deviations from the expected output. Consider potential edge cases, optimization opportunities, or alternative approaches based on insights from both the @@Canonical Solution and @@Execution Result. Offer guidance in a clear and understandable manner, explaining the rationale behind each suggestion. Refrain from providing actual code solutions, but instead focus on conceptual modifications or strategies. Provide constructive feedback to help the programmer improve their coding skills. Remember, your role is to simulate human-like guidance and expertise in programming without directly implementing solutions. Please respond in no more than three sentences. @@Problem {original prompt} @@Existing Code {sanitized code} @@Execution Result {execution result} @@Canonical Solution {canonical solution} @@Guidance
Prompt for GPT models mimicking human feedback without canonical solution You are tasked with providing guidance to a programmer who has drafted a code for a programming problem. Your role is to mimic human-like responses and offer suggestions for modifying the code based on the observed execution results. You should refrain from directly writing code. Begin by thoroughly examining the existing code and its functionality. Analyze the @@Execution Result obtained from running the @@Existing Code. Identify any errors, unexpected behavior, or deviations from the expected output. Consider potential edge cases, optimization opportunities, or alternative approaches based on insights from the @@Execution Result. Offer guidance in a clear and understandable manner, explaining the rationale behind each suggestion. Refrain from providing actual code solutions, but instead focus on conceptual modifications or strategies. Provide constructive feedback to help the programmer improve their coding skills. Remember, your role is to simulate human-like guidance and expertise in programming without directly implementing solutions. Please respond in no more than three sentences. @@Problem {original prompt} @@Existing Code {sanitized code} @@Execution Result {execution result} @@Guidance